# Provably Finding a Hidden Dense Submatrix among Many Planted Dense Submatrices via Convex Programming

### v1

@@ 3.5.4 Completing the Bound on $\boldsymbol{W}$
$\leq\gamma~c_{R}\max\left\{\max_{j}\sqrt{\tilde{\sigma}_{jj}^{2}n_{j}^{*}\log N},\sqrt{\frac{n_{1}}{m_{1}}}\log N\right\}$
@@
$=\frac{1}{\sqrt{m_{1}n_{1}}}+\mathcal{O}(\gamma)$
by our choice of $\tau$ and the assumption that $(\log N)^{3/2}=\mathcal{O}(n_{\min})$.
It follows that $\|\boldsymbol{W}\|\leq 1$ if we choose $\gamma\leq c_{\gamma}/\sqrt{m_{1}n_{1}}$ for sufficiently small $c_{\gamma}>0$ if
$\sqrt{\tilde{\sigma}_{*}^{2}N\log N}=\mathcal{O}(\sqrt{m_{1}n_{1}}).$ (51)
Indeed, (51) implies that
$\frac{1}{\sqrt{m_{1}n_{1}}}\sqrt{\tilde{\sigma}_{jj}^{2}n_{j}^{*}\log N}\leq\frac{1}{\sqrt{m_{1}n_{1}}}\sqrt{\tilde{\sigma}_{*}^{2}N\log N}=\mathcal{O}(1)$
for all $j$, since $n_{j}^{*}\leq N$ and $\tilde{\sigma}_{jj}^{2}\leq\tilde{\sigma}_{*}^{2}$. Moreover,
$\frac{1}{\sqrt{m_{1}n_{1}}}\sqrt{\frac{n_{1}}{m_{1}}}\log N=\frac{\log N}{m_{1}}=\mathcal{O}(1).$
Substituting these bounds in (50) shows that
$\|\boldsymbol{W}\|=\mathcal{O}(c_{\gamma})$
w.h.p. and, hence, choosing sufficiently small $c_{\gamma}$ ensures that $\|\boldsymbol{W}\|<1$ w.h.p.
@@ 5 Numerical Experiments
5.1 An Algorithm for Solving (3)
@@
$\Omega_{Q}$ $:=\{\boldsymbol{Q}\in\mathbb{R}^{M\times N}:P_{\tilde{N}}(\boldsymbol{Q})=\boldsymbol{0}\},$
@@
$\Omega_{Z}$ $:=\{\boldsymbol{Z}\in\mathbb{R}^{M\times N}:0\leq Z_{ij}\leq 1\,\forall(i,j)\in\boldsymbol{v}\times\boldsymbol{v}\}.$
Here, $\mathbb{I}_{S}:\mathbb{R}^{M\times N}\to\{0,+\infty\}$ represents the indicator function of the set $\boldsymbol{S}\subseteq\mathbb{R}^{M\times N}$, where $\mathbb{I}_{S}(\boldsymbol{X})=0$ if $\boldsymbol{X}\in\boldsymbol{S}$, and $\infty$ otherwise.
@@ 5.2.1 Experiment 1
This construction ensures that the $(U_{1},U_{1})$ block is the densest $(m,m)$-submatrix in expectation. Moreover, Conditions (4a) and (4b) hold by our choice of $m$ and $M$. If the gap $q-p_{11}$ is sufficiently large then (4c) is satisfied and we should expect that the matrix representation of $(U_{1},U_{1})$ is the unique optimal solution for (3). When this gap is small, i.e., $q$ is close to $1/2$, then we may not have recovery w.h.p. because the sufficient conditions given by Theorem 2.1 are not satisfied.
This is illustrated by the examples given in Figure 2. Each of the matrices $\boldsymbol{A}$ have a densest block indexed by $(U_{1},U_{1})=([m],[m])$ for $m=100$. The remaining four diagonal blocks are less dense and the off-diagonal blocks are relatively sparse. We can create an undirected graph $G$ with adjacency matrix equal to each sampled matrix $\boldsymbol{A}$. These graphs have densest $m$-subgraphs indexed by $U_{1}$ in expectation. When the gap between $q$ and $p_{22}$, the density of the next densest diagonal block, is small, we have two planted $(m,m)$-submatrices with very similar densities. We cannot expect to recover the densest $m$-subgraph or densest $(m,m)$-submatrix in this case (see Figs. 2(c) and 2(d)). However, when the gap is larger, the corresponding graph $G$ consists of a single dense $m$-subgraph obscured by sparse noise in the form of diversionary edges and nodes. In this case, we should expect to identify the planted dense submatrix and subgraph from the solution of (3) (see Figs. 2(a) and Figs. 2(b)).
@@
The results of this experiment are summarized in Figure 4(a). The observed recovery rates closely match that predicted by Theorem 2.1. Indeed, when $q$ and $m$ satisfy (4c), the solution computed by Alg. 1 is exactly the matrix representation of the planted submatrix in almost all trials.
@@ 5.2.2 Experiment 2
We next consider problem instances of the densest $(m,m)$-submatrix problem constructed as follows for each $(q,m)$-pair in
@@
By construction, $(U_{1},U_{1})$ indexes the densest $(m,m)$-submatrix and $U_{1}$-subgraph of the corresponding undirected graph $G$ when $m>M/2$. If $m<M/2$ then the matrix contains several $(m,m)$-submatrices as dense as that indexed by $(U_{1},U_{1})$; in this case, it is impossible to guarantee that we will recover the matrix representation of $(U_{1},U_{1})$ as the optimal solution of (3); see Figs. 3(g) and 3(h) for an example of such a matrix and graph. On the other hand, the sufficient conditions for recovery given by Theorem 2.1 and (4c), in particular, may not hold if
@@
In the first case, the density in the diagonal blocks is close to that of the off-diagonal blocks and we should not expect to distinguish between them. In the second, the sizes of the two diagonal blocks are close and, again, we should not expect to recover $(U_{1},U_{1})$ as the index set of the densest $(m,m)$-submatrix. However, the sufficient conditions given by Theorem 2.1 suggest that we will if have perfect recovery of $(U_{1},U_{1})$ from the optimal solution of (3) w.h.p. if $m$ and $q$ are sufficiently large (see Figs. 3(a) and 3(b)).
@@
The results of this experiment are summarized in Figure 4(b). As in Section 5.2.1, we have empirical recovery rates that match those predicted by Theorem 2.1. Again, when $q$ and $m$ satisfy (4c) we have perfect recovery of $\boldsymbol{X}_{0}$ in almost all trials. The reported recovery rates are perhaps overly pessimistic in this case. We record recovery of the submatrix indexed by $(U_{1},U_{1})$. When Mm>>M-m>>, this submatrix is not the unique densest $(m,m)$-submatrix in expectation. In this case, the solution $\boldsymbol{X}$ is a convex combination of several $(m,m)$-submatrix of similar densities. We will see in the next section that we can identify the submatrices comprising these solutions using simple rounding techniques in certain cases.
@@
Figure 4: Recovery counts for Experiments 1 and 2. Darker squares indicate more recoveries out of 10 trials for each $(q,m)$-pair.
@@ 5.3.1 Benchmarking Networks
6 Maximum Cliques of the A Song of Ice and Fire Series
We conclude by applying Alg. 1 to identify the largest clique in each of the 5 novels in George R.R. Martin’s A Song of Ice and Fire (ASOIAF) saga using the character interaction networks considered in Beveridge and Shan (2016); Each network, one for each book in series, has nodes corresponding to characters and edge weights corresponding to the number of times two character names appear within 15 words of one another, which models the number of interactions between each pair of characters in the book. A detailed discussion of the properties of these interaction networks, as well as all network data are available at 37. A similar analysis of the television adaptation of ASOIAF, HBO’s Game of Thrones, is presented in Beveridge and Chemers (2018).
@@ 6 Maximum Cliques of the A Song of Ice and Fire Series
The maximum clique in each interaction network represents the largest collection of fully connected characters in the corresponding book. Here, the membership of these cliques reflects the dynamics of the series of novels. In the first book, we observe a single large community primarily comprised of the Baratheon, Stark, and Lannister families. As the saga progresses, the action of the novels scatters these characters across many disparate locations. This corresponds to interaction networks consisting of much smaller cliques and clusters. For each book and interaction network, we successfully identified the largest of these cliques, but further analysis is required to find all of the underlying communities. This analysis aligns with the more detailed analysis of the network dynamics of the ASOIAF series found in 10; 9; 37.

### v2

@@ 3.5.4 Completing the Bound on $\boldsymbol{W}$
$\leq\gamma~c_{R}\left(\max\left\{\sqrt{\sigma_{11}^{2}n_{1}^{*}\log N},\sqrt{\frac{n_{1}^{*}}{n_{1*}}}\log N\right\}\right)$
@@
$=\frac{1}{\sqrt{m_{1}n_{1}}}+\mathcal{O}(\gamma)=\mathcal{O}(\gamma)$
if we choose
$\gamma=\mathcal{O}\left(\frac{1}{(p_{11}-p^{*})\sqrt{m_{1}n_{1}}}\right)$
by our choice of $\tau$ and the assumption that $(\log N)^{3/2}=\mathcal{O}(n_{\min})$ in (4b).
It follows that $\|\boldsymbol{W}\|\leq 1$ and (31), (34), and (36), which ensure nonnegativity of $\boldsymbol{\Lambda}$ and $\boldsymbol{\Xi}$, are satisfied with high probability if we choose $\gamma\leq c_{\gamma}/((p_{11}-p^{*})\sqrt{m_{1}n_{1}})$ for sufficiently small $c_{\gamma}>0$ and
$p_{11}-p_{rs}\geq c_{1}\max\left\{\sqrt{\frac{\tilde{\sigma}^{2}N^{*}\log N^{*}}{m_{1}n_{1}}},\sqrt{\frac{\sigma_{11}^{2}\log N^{*}}{\min\{m_{1},n_{1}\}}},\sqrt{\frac{\max\{\tilde{\sigma}^{2},1\}(\log N^{*})^{3}}{\min\{m_{1},n_{1}\}}}\right\}$
for all $(r,s)\neq(1,1).$ This completes the proof.
@@ 5 Numerical Experiments
5.1 An Algorithm Based on the Alternating Direction Method of Multipliers
@@
$\Omega_{Q}$ $:=\{\boldsymbol{Q}\in\mathbb{R}^{M\times N}:P_{\Omega}(\boldsymbol{Q})=\boldsymbol{0}\},$
@@
$\Omega_{Z}$ $:=\{\boldsymbol{Z}\in\mathbb{R}^{M\times N}:0\leq Z_{ij}\leq 1\,\forall(i,j)\}.$
Here, $\mathbb{I}_{\mathcal{S}}:\mathbb{R}^{M\times N}\to\{0,+\infty\}$ represents the indicator function of the set $\mathcal{S}\subseteq\mathbb{R}^{M\times N}$, where $\mathbb{I}_{\mathcal{S}}(\boldsymbol{X})=0$ if $\boldsymbol{X}\in\mathcal{S}$, and $\infty$ otherwise.
@@
Remark 3 (Computational Cost of Alg. 1).
We need $\mathcal{O}(MN)$ floating point operations (flops) to update all primal and dual variables except $\boldsymbol{X}$ during each iteration of Alg. 1. Indeed, each of these variable updates reduces to entry-wise operations on $M\times N$ real matrices. Updating $\boldsymbol{X}$ is the most computationally expensive step of each iteration and requires the calculation of the singular value decomposition of a dense matrix. This requires $\mathcal{O}(\min\{M,N\}MN)$ flops per iteration. Thus, each iteration of Alg. 1 uses $\mathcal{O}(\min\{M,N\}MN)$ flops.
@@ 5.2.1 Experiment 1
This construction ensures that the $(U_{1},U_{1})$ block is the densest $m\times m$-submatrix in expectation. Moreover, Conditions (4a) and (4b) hold by our choice of $m$ and $M$. If the gap $q-p_{11}$ is sufficiently large then (4c) is satisfied and we should expect that the matrix representation of $(U_{1},U_{1})$ is the unique optimal solution for (3). When this gap is small, i.e., $q$ is close to $1/2$, then we may not have recovery w.h.p. because the sufficient conditions given by Theorem 2.1 are not satisfied.
This is illustrated by the examples given in Figure 2. Each of the matrices $\boldsymbol{A}$ have a densest block indexed by $(U_{1},U_{1})=([m],[m])$ for $m=100$. The remaining four diagonal blocks are less dense and the off-diagonal blocks are relatively sparse. We can create an undirected graph $G$ with adjacency matrix equal to each sampled matrix $\boldsymbol{A}$. These graphs have densest $m$-subgraphs indexed by $U_{1}$ in expectation. When the gap between $q$ and $p_{22}$, the density of the next densest diagonal block, is small, we have two planted $(m,m)$-submatrices with very similar densities. We cannot expect to recover the densest $m$-subgraph or densest $m\times m$-submatrix in this case (see Figs. 2(c) and 2(d)). However, when the gap is larger, the corresponding graph $G$ consists of a single dense $m$-subgraph obscured by sparse noise in the form of diversionary edges and nodes. In this case, we should expect to identify the planted dense submatrix and subgraph from the solution of (3) (see Figs. 2(a) and Figs. 2(b)).
@@
The results of this experiment are summarized in Figure 4(a). Figure 4(a) includes the theoretical phase transition curve
$q=0.25+\max\left\{\sqrt{\frac{M\log M}{3m}},\frac{\log M}{m}\right\}$
given by (4c) with constant $c_{1}=1$. The observed recovery rates closely match that predicted by Theorem 2.1. Indeed, when $q$ and $m$ satisfy (4c), the solution computed by Alg. 1 is exactly the matrix representation of the planted submatrix in all trials.
@@ 5.2.2 Experiment 2
We next consider problem instances of the densest $m\times m$-submatrix problem constructed as follows for each $(q,m)$-pair in
@@
By construction, $(U_{1},U_{1})$ indexes the densest $m\times m$-submatrix and $U_{1}$-subgraph of the corresponding undirected graph $G$ when $m>M/2$. If $m<M/2$ then the matrix contains several $(m,m)$-submatrices as dense as that indexed by $(U_{1},U_{1})$; in this case, it is impossible to guarantee that we will recover the matrix representation of $(U_{1},U_{1})$ as the optimal solution of (3); see Figs. 3(g) and 3(h) for an example of such a matrix and graph. On the other hand, the sufficient conditions for recovery given by Theorem 2.1 and (4c), in particular, may not hold if
@@
In the first case, the density in the diagonal blocks is close to that of the off-diagonal blocks and we should not expect to distinguish between them. In the second, the sizes of the two diagonal blocks are close and, again, we should not expect to recover $(U_{1},U_{1})$ as the index set of the densest $m\times m$-submatrix. However, the sufficient conditions given by Theorem 2.1 suggest that we will if have perfect recovery of $(U_{1},U_{1})$ from the optimal solution of (3) w.h.p. if $m$ and $q$ are sufficiently large (see Figs. 3(a) and 3(b)).
@@
The results of this experiment are summarized in Figure 4(b); we plot the theoretical phase transition curve
$q=0.25+6\,\max\left\{\sqrt{\frac{M\log M}{3m}},\frac{\log M}{m}\right\}$
given by (4c) with constant $c_{1}=6$. As in Section 5.2.1, we have empirical recovery rates that match those predicted by Theorem 2.1. Again, when $q$ and $m$ satisfy (4c) we have perfect recovery of $\boldsymbol{X}_{0}$ in almost all trials. The reported recovery rates are perhaps overly pessimistic in this case. We record recovery of the submatrix indexed by $(U_{1},U_{1})$. When $M>>m$, this submatrix is not the unique densest $m\times m$-submatrix in expectation. In this case, the solution $\boldsymbol{X}$ is a convex combination of several $m\times m$-submatrices of similar densities. We will see in the next section that we can identify the submatrices comprising these solutions using simple rounding techniques in certain cases.
@@
Figure 4: Recovery counts for Experiments 1 and 2. Darker squares indicate more recoveries out of 10 trials for each $(q,m)$-pair. The dashed curve indicates the phase transition curve to perfect recovery given by the right-hand side of (4c)
@@ 5.3.1 Benchmarking Networks
5.4 Maximum Cliques of the A Song of Ice and Fire Series
We conclude by applying Alg. 1 to identify the largest clique in each of the 5 novels in George R.R. Martin’s A Song of Ice and Fire (ASOIAF) saga using the character interaction networks considered in Beveridge and Shan (2016); Each network, one for each book in series, has nodes corresponding to characters and edge weights corresponding to the number of times two character names appear within 15 words of one another, which models the number of interactions between each pair of characters in the book. A detailed discussion of the properties of these interaction networks, as well as all network data are available at 39. A similar analysis of the television adaptation of ASOIAF, HBO’s Game of Thrones, is presented in Beveridge and Chemers (2018).
@@ 5.4 Maximum Cliques of the A Song of Ice and Fire Series
The maximum clique in each interaction network represents the largest collection of fully connected characters in the corresponding book. Here, the membership of these cliques reflects the dynamics of the series of novels. In the first book, we observe a single large community primarily comprised of the Baratheon, Stark, and Lannister families. As the saga progresses, the action of the novels scatters these characters across many disparate locations. This corresponds to interaction networks consisting of much smaller cliques and clusters. For each book and interaction network, we successfully identified the largest of these cliques, but further analysis is required to find all of the underlying communities. This analysis aligns with the more detailed analysis of the network dynamics of the ASOIAF series found in 11; 10; 39.
@@
5.5 Choice of Regularization Parameter
The sufficient conditions for perfect recovery given in Section 2 hold for choice of the regularization parameter $\gamma$ within a certain range. The endpoints of this interval are proportional to $1/\sqrt{m_{1}n_{1}}$ and depend on the probabilities $p_{11}$ and $p^{*}$. In practice, we may not know these probabilities.
In our experiments involving synthetic data described in Sections 5.2.1 and 5.2.2, we use the value of $p_{11}$ and $p^{*}$ used to construct our random matrices. Specifically, we
$\gamma=\frac{6}{(p_{11}-p^{*})\sqrt{mn}}=\frac{6}{\sqrt{mn}(q-p)},$
since the numerator $6$ is in the range of coefficients satisfying (5). In our analysis of real-world networks in Section 5.3, we try to identify the densest submatrix with area equal to the square of the size of the maximum clique of each network. Here, we can assume that $p_{11}=1$ since the block indexed by the maximum clique is fully dense. On the other hand, we use the rough estimate $p^{*}\approx 1/2$ in our formula for $\gamma$:
$\gamma=\frac{12}{m}=\frac{6}{m/2}\approx\frac{6}{m(p_{11}-p^{*}}.$
In practice, if either $p_{11}$ or $p^{*}$ are not known, we suggest using simple estimation procedures combined with cross-validation to choose $\gamma$. For example, $p_{11}$ can be approximated by normalizing the maximum row or column sum of $\boldsymbol{A}$, while $p^{*}$ can be approximated using the density of the graph or a similar function of row and column sums of $\boldsymbol{A}$. We can then use cross-validation to select $\gamma$ from an interval centered at the choice of $\gamma$ given by these approximations.
5.5.1 Robustness to Choice of Regularization Parameter
To test the robustness of the recovery guarantees given by Theorem 2.1 and Theorem 2.2 to the choice of the regularization parameter $\gamma$, we performed the following experiment.
  • For each $\gamma$ in a set of potential values of the regularization parameter, we solve (3) using Alg. 1.
  • For each choice of $\gamma$, we stop Alg. 1 after a $10^{-4}$-suboptimal solution is found or $2000$ iterations have been performed.
  • We then compare the recovered solution $\boldsymbol{X}$ with the indicator matrix $\boldsymbol{X}_{0}$ for the known densest submatrix before and after rounding each entry of $\boldsymbol{X}_{0}$ to the nearest integer.
5.5.1.1  Synthetic Data
We first sample $10$ matrices from the planted submatrix model as follows:
  • We set $M,N=500$ and partition $[M]$ and $[N]$ into two blocks with $m_{1},n_{1}=400$ and $m_{2},n_{2}=100$.
  • We set $p_{11}=0.75$ and $p_{12}=p_{21}=p_{22}=0.25$.
The sufficient condition given in Theorem 2.1 suggests that we should have perfect recovery of the hidden dense $m_{1}\times n_{1}$-submatrix for certain choices of $\gamma$ w.h.p. in this case.
We apply Alg. 1 to solve each instance of (3) corresponding to each
$\gamma\in\{0,0.005,0.01,0.015,\dots,0.145,0.15\}.$
This set includes the value used in Sections 5.2.1 and 5.2.2
$\gamma=\frac{12}{\sqrt{mn}}=0.03.$
The result of this experiment is summarized in Figure 8. Here, we observe that recovery error is below the stopping tolerance $10^{-4}$ for all $\gamma$ chosen within the interval $[0.02,0.12]$ with poor recovery error outside this interval. This agrees with the claim that we should have perfect recovery (up to error from early termination) for all $\gamma$ within a certain interval.
(a) Recovery Error
(b) Recovery Error After Rounding
(c) Run-time (s)
Figure 8: Average recovery error with $95\%$-confidence interval before and after rounding for each choice of $\gamma$, as well as average run-time, for $10$ matrices sampled from the planted submatrix model where we should expect perfect recovery for well-chosen $\gamma$.
5.5.1.2  Real-World Networks
We repeated the analysis of Section 5.5.1.1 with two of the networks considered in Sections 5.3 and 5.4:
  • The JAZZ Network with $m=n=30$ and $\gamma\in\{0,0.008,0.016,0.024,\dots,0.392,0.4\}$; and
  • Book 1 of the ASOIAF series with $m=n=18$ and $\gamma\in\{0,0.02,0.04,0.06,\dots,0.98,1\}$.
For each network and choice of $m$ and $\gamma$, we call Alg. 1 and compare the calculated solution with the known maximum clique of size $m$. The recovery error before and after rounding is visualised in Figures 9 and 10. In both cases, we can observe that we have perfect recovery of the maximum clique for all $\gamma$ above a critical value. For example, we have recovery, before and after rounding, if we use
$\gamma=\frac{6}{m(1-\bar{p})},$ (66)
indicated by the dashed vertical line in Figures 9 and 10, where we estimate the unknown parameter $p^{*}$ using the $\bar{p}=\frac{1}{m^{2}}\sum_{i,j}a_{ij}$ is the density of the network (with $\boldsymbol{A}$ denoting the adjacency matrix of the network).
(a) Recovery Error
(b) Recovery Error After Rounding
(c) Run-time (s)
Figure 9: Recovery error before and after rounding, and run-time, for each choice of $\gamma$ for the JAZZ collaboration network. The dashed vertical line denotes the choice of $\gamma$ using density as an estimate of $p^{*}$ according to (66).
(a) Recovery Error
(b) Recovery Error After Rounding
(c) Run-time (s)
Figure 10: Recovery error before and after rounding, and run-time, for each choice of $\gamma$ for the ASOIAF Book 1 interaction network. The dashed vertical line denotes the choice of $\gamma$ using density as an estimate of $p^{*}$ according to (66).
5.5.1.3  The Benefits of Rounding
In each of these experiments, we have perfect recovery after rounding for a wider range of $\gamma$ than without rounding. When the hypothesis of Theorems 2.1 and 2.2 are satisfied the solution at termination of Algorithm 1 is close to the rank-one matrix representation of the densest $m\times n$-submatrix; in this case, rounding will give exactly the rank-one indicator matrix.
When $\gamma$ is chosen outside the range of values given in (5), the rank-one matrix representation may not be optimal for (3). In this case, we still may be able to identify the densest $m\times n$-submatrix using rounding. Figures 11(a)12(b), and 13(b) illustrate this phenomena. In each case, we find an optimal solution $\boldsymbol{X}^{*}$ of the convex relaxation (3) with strictly fractional entries. Rounding the entries of $\boldsymbol{X}^{*}$ to the nearest integer yields the matrix representation of the planted $m\times n$-submatrix in each of these cases; this rank-one binary matrix is suboptimal for (3) with this choice of $\gamma$ but does reveal the solution for the original combinatorial problem. However, if $\gamma$ is chosen too far outside this interval, we obtain optimal solutions for the relaxation (3) which are far from the optimal solution of the original problem, as illustrated by Figures 11(c)12(a), and 13(a). In this case, we cannot extract the densest $m\times n$-submatrix from these optimal solutions of (3) by rounding.
(a) $\gamma=0.005$
(b) $\gamma=0.03$
(c) $\gamma=0.15$
Figure 11: Recovered solutions for a random matrix sampled from the planted submatrix model from Sect. 5.5.1.1 for different choices of $\gamma$.
(a) $\gamma=0.008$
(b) $\gamma=0.048$
(c) $\gamma=0.4$
Figure 12: Recovered solutions for the JAZZ Network for different choices of $\gamma$.
(a) $\gamma=0.02$
(b) $\gamma=0.14$
(c) $\gamma=$0.46
Figure 13: Recovered solutions for the ASOIAF Book 1 Network for different choices of $\gamma$.